40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2347     Entry time: Thu May 16 13:46:46 2019
Author: anchal 
Type: DailyProgress 
Category: TempCtrl 
Subject: Alignment disturbing due to lab temp changes 

As suspected in CTN: 2346, the mode matching of the cavities is deteriorating and eventually alignment is getting screwed due to possible lab temperature fluctuations. I left yesterday with south cavity mode matching to about ~70% but in the morning today, I found that the resonance is completely lost and a higher order mode with vertical fringes is resonant. Same is the case with North one which had shifted to a much higher order with vertical fringes. So clearly, I need to switch back on the vacuum can temperature stabalization.

ELOG V3.1.3-